# Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions

### v2

@@ Appendix C Omitted Proofs from Section 5
With this lemma, we are ready to prove Lemma C.2.
Case I: $Q_{j}=\widehat{Q}_{j}$. By assumption, $Q_{j}=\widehat{Q}_{j}\leq Q_{j}^{*}$. So, for any $t\in T_{j,1}$, invoking Lemma C.5 (1) with $q=\widehat{Q}_{j}$, we conclude that $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ is allocated exactly $\widehat{Q}_{j}$ units. Summing over all rounds in $T_{j,1}$,
@@
Now, suppose $T_{j,0}\neq\emptyset$ and $r_{t}^{*}=\widehat{Q}_{j}$ for some $t\in T_{j,0}$. By definition, $\widehat{Q}_{j}<Q_{j}^{*}$ and $\widehat{Q}_{j}\geq r_{s}^{*}$ for any $s\in T_{j,0}$. So, for any $s\in T_{j,0}$, invoking Lemma C.5 (2) with $r_{t}^{*}=\widehat{Q}_{j}$, we conclude that $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ is allocated at least $r_{s}^{*}$ units. Hence, summing over all rounds, $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ gets at least the value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ over the rounds in $T_{j,0}$. So,
@@
If $T_{j,0}=\emptyset$, the lower bound is trivially true. Combining Eq. 31 and (32), for $Q_{j}=\widehat{Q}_{j}$,
@@
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-})=\frac{\max_{\mathbf{b}^{\prime}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$
Then, maximizing over all bid histories, we get the desired result.
Let $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$. If $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is a $k$-uniform bidding strategy where $k\leq m\leq m^{\prime}$, by definition, $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)\in\mathscr{F}^{\mathcal{H}_{-}}_{m}$ implying $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-})=1\leq\frac{m^{\prime}}{m}$.
Without loss of generality, let the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$ be $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b^{*}_{1},q^{*}_{1}),\dots,(b^{*}_{m^{\prime}},q^{*}_{m^{\prime}})\rangle$ and $Q_{\ell}^{*}=\sum_{j\leq\ell}q_{j}^{*}$ for all $\ell\in[m^{\prime}]$ [^13] A similar analysis also follows if $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is a $k$-uniform strategy where $m\leq k\leq m^{\prime}$. Let $r_{t}^{*}$ be the number of units allocated to $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in round $t$. Recall that $T_{j}$ is the set of rounds when $b_{j}^{*}$ is the least winning bid when $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is submitted (see Eq. 15). So,
@@
0:  $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b^{*}_{1},q^{*}_{1}),\dots,(b^{*}_{m^{\prime}},q^{*}_{m^{\prime}})\rangle$. The collection, $\mathscr{C}$, of all subsets $\mathcal{S}\subseteq[m^{\prime}]$ such that $|S|=m$. Initialize $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}=\emptyset$.
1:  for $\mathcal{S}\in\mathscr{C}$ do
2:     Suppose $\mathcal{S}=\{k_{1},\dots,k_{m}\}$ such that ${k_{1}}<{k_{2}}<\dots<{k_{m}}$.
3:     Construct a $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ where
$b_{j}=b_{k_{j}}^{*}\quad\text{and}\quad q_{j}=Q_{k_{j}}^{*}-Q_{k_{j-1}}^{*},~{}$ $\forall j\in[m]\,.$ (39)
4:     $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\leftarrow\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\cup\{\mathbf{b}\}$
5:  end forReturn: The class of bidding strategies, $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$.
@@
Recall that $T_{j}$s, the set of rounds in which $b_{j}^{*}$ is the least winning bid, form a partition of $[T]$. So,
$\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{t=1}^{m^{\prime}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})=\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})+\sum_{j\in[m^{\prime}]\setminus\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})\geq\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (42)
@@
$V(\mathbf{b};\bm{\beta}_{-}^{t})$ =(41)${{=}}\max_{\ell\in\mathcal{S}}V((b_{\ell}^{*},Q_{\ell}^{*});\bm{\beta}_{-}^{t})\geq V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})\implies\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})\geq\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})$ ) (43)
@@
(44)${{\geq}}\max_{\mathcal{S}\subseteq[m^{\prime}]:|\mathcal{S}|=m}\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})$ )
@@
Maximizing over all bid histories, we get that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}\leq\frac{m^{\prime}}{m}$.
Now, we prove that the result also holds for safe bidding strategies. Formally, we show that for $m,m^{\prime}\in\mathbb{N}$ and $m^{\prime}\geq m$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}\leq\frac{m^{\prime}}{m}$. To this end, we consider the quantity $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ analogous to $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-})$, upper bound it and maximize over all bid histories to get the desired result.
@@
For any subset $\mathcal{S}\subseteq[m^{\prime}]$ such that $|S|=m$, suppose $\mathbf{b}$ is the bidding strategy obtained from $\mathcal{S}$. Then, $\mathbf{b}$ is a safe bidding strategy. As $\mathcal{S}$ was chosen arbitrarily, $\mathscr{S}_{m}(\mathcal{H}_{-})\subseteq\mathscr{S}_{m}$.
@@ Appendix D Tight Lower Bounds for Results in Section 5
D.1.  Tight Lower Bound for Theorem 5.3 (For $m=1$)
In this section, we construct a bid history, $\mathcal{H}_{-}$ and valuation curve $\mathbf{v}$ (equivalently the $\mathscr{U}_{m}^{\star}$) for which the upper bound on $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}$, presented in Theorem 5.3, is tight. Recall that for any choice of $\mathcal{H}_{-}$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\leq 2-\theta_{\mathcal{H}_{-}}$, where $\theta_{\mathcal{H}_{-}}\leq\max_{j\in[m]}({W}_{\widehat{Q}j}/{W}_{Q_{j}^{*}})$. To minimize $\theta_{\mathcal{H}_{-}}$, we choose a valuation curve that is very weakly decreasing. We then set the competing bids such that $\widehat{Q}_{j}\ll Q_{j}^{*}$ for all $j\in[m]$. Finally, we determine the values of $|T_{j,0}|$ and $|T_{j,1}|$ for which the upper bound is tight. See the definition of these quantities in Section C.1, where the first part of the theorem is proven.
We present the case when $m=1$ below. The case for $m\geq 2$ is deferred to Section D.2. Although the main idea for both the cases are the same, for $m\geq 2$, the construction is more involved, and hence presented separately. Formally, for any $\delta\in(0,1/2]$, we design a bid history, $\mathcal{H}_{-}$, and valuation vector, $\mathbf{v}$ for which $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{1},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq 2-\delta$.
Let $M=2\left\lceil\frac{1}{\delta}\right\rceil$. Suppose $\mathbf{v}=[1,v,\cdots,v]\in\mathbb{R}^{M}$, target RoI $\gamma=0$ and $v=1-4\epsilon$ where $\epsilon=\frac{\delta-1/M}{4(1-1/M)}<\frac{\delta}{4}$. Observe that $\epsilon\in(0,1/8)$ as $\delta\leq 1/2$. Let $T=M$ and $K=M+1$. The bid history is defined as:
$\bm{\beta}_{-,t}^{-(j)}=\begin{cases}1-\epsilon,~{}&\text{ if $t\leq M-1$ and $j=1$}\\ C,~{}&\text{ if $t\leq M-1$ and $2\leq j\leq K$}\\ \epsilon,~{}&\text{ if $t=M$ and $1\leq j\leq K$}\\ \end{cases}\,,$
where $C\gg w_{1}$. The bid history is presented in Table 1.
Table 1: Bid history, $\mathcal{H}_{-}$, that achieves the $2$-approximation asymptotically for $m=1$. Here $C\gg w_{1}$ and $\epsilon>0$ is a small real number.
$t=1$ $t=2$cdots $t=M-1$ $t=M$
$C$ $C$cdots $C$ $\epsilon$
$\vdots$ $\vdots$ $\cdots$ $\vdots$ $\vdots$
$C$ $C$cdots $C$vdots
$1-\epsilon$ $1-\epsilon$cdots $1-\epsilon$ $\epsilon$
Notice that the constructed $\mathcal{H}_{-}$ contains $M-1$ rounds with high competition (where the bidder can acquire at most 1 unit), while there is one round with minimal competition, allowing the bidder to obtain any desired number of units. The 1-uniform safe bidding strategies of the form $(w_{q},q)$ do not perform well universally on all rounds because with increase in $q$, despite the increasing demand (i.e., $q$), the bid value $w_{q}$ decreases, thereby reducing the likelihood of acquiring any units.
Notations. For a given bid history $\mathcal{H}_{-}$, recall that for any $m\in\mathbb{N}$, $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$. We define $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)=\sum_{t=1}^{T}V(\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})$, the value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ over the bid history $\mathcal{H}_{-}$. Similarly, $\mathbf{b}_{m}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right):=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ and $V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)=\sum_{t=1}^{T}V(\mathbf{b}_{m}^{\mathsf{SAFE}}\left(\mathcal{H}_{-}\right);\bm{\beta}_{-}^{t})$.
Computing the optimal strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{1}$. We observe that no bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{1}$ can obtain more than 1 unit in the first $M-1$ rounds and $M$ units in the final round.
We claim that $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=(1,M)$. Observe that $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is allocated 1 unit in each of the first $M-1$ rounds and $M$ units in the final round. Hence, it is allocated the maximum number of units possible. To verify that $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is feasible for the bid history, notice that for $\forall t\in[M-1]$, $p(\bm{\beta}^{t})=1\leq w_{1}=1$. For round $t=M$, as $K>M$,
$p(\bm{\beta}^{t})=\epsilon\leq 1-4\epsilon<w_{M}=\mu\cdot 1+(1-\mu)\cdot(1-4\epsilon),\quad\text{where}\quad\mu=\frac{1}{M},$
implying that RoI constraint is satisfied by $\mathbf{b}_{1}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in round $t=T$. So, $V^{\mathsf{OPT}}_{1}\left(\mathcal{H}_{-}\right)=M+v(M-1)$.
Computing the optimal strategy in $\mathscr{U}_{1}^{\star}$. Here, we argue that the optimal strategy in $\mathscr{U}_{1}^{\star}$ is $(w_{1},1)=(1,1)$. Observe that the strategy $(1,1)$ is allocated 1 unit in each round. Hence, it obtains a total value of $M$.
As $\bm{\beta}_{-,t}^{-(1)}=1-\epsilon>1-2\epsilon=w_{2}$, for $t=1,2,\dots,M-1$, bidding $\left(w_{q},q\right)$ for $q\geq 2$ does not get any value in the first $M-1$ rounds. Bidding $\left(w_{q},q\right)$ gets exactly $q$ units in round $M$ as $w_{q}>\epsilon$ for any $q\geq 2$. So, for $2\leq q\leq M$, the total value obtained by bidding $(w_{q},q)$ is $1+(q-1)v<q$. Hence, $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)=M$, which is the value obtained by the safe strategy $(w_{1},1)=(1,1)$. So,
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{1};\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{V^{\mathsf{OPT}}_{1}\left(\mathcal{H}_{-}\right)}{V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)}=\frac{M+v(M-1)}{M}=\frac{2M-1-4\epsilon(M-1)}{M}=2-\delta\,.$
D.2.  Tight Lower Bound for Theorem 5.3 (For $m\geq 2$)
@@
where the second inequality follows because $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}\leq m$ (by Theorem 5.4). So, instead of computing $V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)$ directly which requires solving a DP, we obtain $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$ and show that the bound is tight.
D.2.1.  Construction of $\mathcal{H}_{-}$.
@@
Table 2: Example of bid history, $\mathcal{H}_{-}$, achieving 2-approximation asymptotically for $m=2$. Each round in the same partition has identical competing bid profile. Total number of units in each auction is $K=N^{3}+1$.
@@
D.2.2.  Computing $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$.
We make the following claim about the optimal $m$-uniform bidding strategy for the constructed $\mathcal{H}_{-}$.
@@
D.2.3.  Computing $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$

### v4

@@ Appendix C Omitted Proofs from Section 5
With this lemma, we are ready to prove Lemma 5.7.
Case I: $Q_{j}=\widehat{Q}_{j}<Q_{j}^{*}$. For any $t\in T_{j,1}$, invoking Lemma C.3 (1) with $q=\widehat{Q}_{j}$, we conclude that $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ is allocated exactly $\widehat{Q}_{j}$ units. Summing over all rounds in $T_{j,1}$,
@@
As $\widehat{Q}_{j}<Q_{j}^{*}$, $T_{j,0}\neq\emptyset$. Let $r_{t}^{*}=\widehat{Q}_{j}$ for some $t\in T_{j,0}$. By definition, $\widehat{Q}_{j}<Q_{j}^{*}$ and $\widehat{Q}_{j}\geq r_{s}^{*}$ for any $s\in T_{j,0}$. So, for any $s\in T_{j,0}$, invoking Lemma C.3 (2) with $r_{t}^{*}=\widehat{Q}_{j}$, we conclude that $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ is allocated at least $r_{s}^{*}$ units. Hence, summing over all rounds, $(w_{\widehat{Q}_{j}},\widehat{Q}_{j})$ gets at least the value obtained by $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ over the rounds in $T_{j,0}$. So,
@@
Combining Eq. 40 and (41), for $Q_{j}=\widehat{Q}_{j}$,
@@
C.3. Proof of Lemma 5.8
Let $(\mathcal{S}_{0}^{*},\widehat{\mathcal{S}}_{0})=\mathop{\mathrm{argmax}}_{(\mathcal{S}^{*},\widehat{\mathcal{S}})}\Big\{\sum_{j\in\mathcal{S}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)\Big\}$. Then,
$=\sum_{j\in\mathcal{S}_{0}^{*}}\Big(V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big)\,.$ (46)
Consider a partition of $[m]$ that differs from the maximizing partition $(\mathcal{S}_{0}^{*},\widehat{\mathcal{S}}_{0})$ by exactly one element, i.e., for any $a\in\widehat{\mathcal{S}}_{0}$, consider the following partition: $(\mathcal{S}_{0}^{*}\cup\{a\},\widehat{\mathcal{S}}_{0}\setminus\{a\})$. By definition,
$\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)$ $\geq\sum_{j\in\mathcal{S}_{0}^{*}\cup\{a\}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}\setminus\{a\}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)$
Now, for any $b\in\mathcal{S}^{*}_{0}$, consider the following partition: $(\mathcal{S}_{0}^{*}\setminus\{b\},\widehat{\mathcal{S}}_{0}\cup\{b\})$. By definition,
$\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)$ $\geq\sum_{j\in\mathcal{S}_{0}^{*}\setminus\{b\}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}\cup\{b\}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)$
Plugging in the values,
$\frac{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}{\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)}$
$=\frac{\sum_{j\in\mathcal{S}_{0}^{*}}\Big(V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{Q_{j}^{*}}|T_{j,1}|\Big)}{\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)}$
[…]
$=\frac{2\Big\{\sum_{j\in\mathcal{S}_{0}^{*}}{W}_{Q_{j}^{*}}|T_{j,1}|+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)\Big\}-\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)}$
$=2-\theta_{\mathcal{H}_{-}}\,$,
where
$0<\theta_{\mathcal{H}_{-}}$ $=\frac{\sum_{j=1}^{m}{W}_{\widehat{Q}_{j}}|T_{j,1}|}{\sum_{j\in\mathcal{S}_{0}^{*}}\Big({W}_{Q_{j}^{*}}|T_{j,1}|\Big)+\sum_{j\in\widehat{\mathcal{S}}_{0}}\Big(V_{j,0}|T_{j,0}|+{W}_{\widehat{Q}_{j}}|T_{j,1}|\Big)}$
[…]
and $(a)$ follows because for positive $a_{1},\dots,a_{m}$ and $b_{1},\dots,b_{m}$, $\frac{\sum_{j=1}^{m}a_{j}}{\sum_{j=1}^{m}b_{j}}\leq\max_{j\in[m]}\frac{a_{j}}{b_{j}}$.
@@
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-},\mathbf{v})=\frac{\max_{\mathbf{b}^{\prime}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b}^{\prime};\bm{\beta}_{-}^{t})}{\max_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})}\,.$
Then, maximizing over all bid histories and valuation vectors, we get the desired result.
Let $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\mathop{\mathrm{argmax}}_{\mathbf{b}\in\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$. If $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is a $k$-uniform bidding strategy where $k\leq m\leq m^{\prime}$, by definition, $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)\in\mathscr{F}^{\mathcal{H}_{-}}_{m}$ implying $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-},\mathbf{v})=1\leq\frac{m^{\prime}}{m}$.
Without loss of generality, let the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$ be
$\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)=\langle(b^{*}_{1},q^{*}_{1}),\dots,(b^{*}_{m^{\prime}},q^{*}_{m^{\prime}})\rangle$
and $Q_{\ell}^{*}=\sum_{j\leq\ell}q_{j}^{*}$ for all $\ell\in[m^{\prime}]$. [^23] A similar analysis also follows if $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is a $k$-uniform strategy where $m\leq k\leq m^{\prime}$. Let $r_{t}^{*}$ be the number of units allocated to $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ in round $t$. Recall that $T_{j}$ is the set of rounds when $b_{j}^{*}$ is the least winning bid when $\mathbf{b}_{m^{\prime}}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$ is submitted (see Eq. 11). So,
@@
[…]
2:Initialize $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\leftarrow\emptyset$.
3:for $\mathcal{S}\in\mathscr{C}$ do
4:  Suppose $\mathcal{S}=\{k_{1},\dots,k_{m}\}$ such that ${k_{1}}<{k_{2}}<\dots<{k_{m}}$.
5:  Construct a $m$-uniform bidding strategy $\mathbf{b}=\langle(b_{1},q_{1}),\dots,(b_{m},q_{m})\rangle$ where
$b_{j}=b_{k_{j}}^{*}\qquad\text{and}\qquad q_{j}=Q_{k_{j}}^{*}-Q_{k_{j-1}}^{*},$~ $\forall j\in[m]\,.$ (51)
6:  $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\leftarrow\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}\cup\{\mathbf{b}\}$
7:end for
8:return The class of bidding strategies, $\widetilde{\mathscr{G}}^{\mathcal{H}_{-}}_{m}$.
@@
Recall that $T_{j}$s, the set of rounds in which $b_{j}^{*}$ is the least winning bid, partition the set $[T]$. So,
$\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $=\sum_{t=1}^{m^{\prime}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$
$=\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})+\sum_{j\in[m^{\prime}]\setminus\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (54)
$\geq\sum_{j\in\mathcal{S}}\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (55)
@@
[…]
$\implies\sum_{t\in T_{j}}V(\mathbf{b};\bm{\beta}_{-}^{t})$ $\geq\sum_{t\in T_{j}}V((b_{j}^{*},Q_{j}^{*});\bm{\beta}_{-}^{t})$ (56)
@@
[…]
@@
Maximizing over all bid histories and valuation vectors, we get that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}\leq\frac{m^{\prime}}{m}$.
Now, we prove that the result also holds for safe bidding strategies. Formally, we show that for $m,m^{\prime}\in\mathbb{N}$ and $m^{\prime}\geq m$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}\leq\frac{m^{\prime}}{m}$. To this end, we consider the quantity $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ analogous to $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{F}^{\mathcal{H}_{-}}_{m}}(\mathcal{H}_{-},\mathbf{v})$, upper bound it and maximize over all bid histories and valuation vectors to get the desired result.
@@
For any subset $\mathcal{S}\subseteq[m^{\prime}]$ such that $|\mathcal{S}|=m$, suppose $\mathbf{b}$ is the bidding strategy obtained from $\mathcal{S}$ by applying Algorithm 2. Then, $\mathbf{b}$ is a safe bidding strategy. As $\mathcal{S}$ was chosen arbitrarily, $\mathscr{S}_{m}(\mathcal{H}_{-})\subseteq\mathscr{S}_{m}$.
@@ Appendix D Tight Lower Bounds for Results in Section 5
D.1. Tight Lower Bound for Theorem 5.3 (For $m\geq 2$)
@@
where the second inequality follows because $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}\leq m$ (by Theorem 5.4). Surprisingly, we show that instead of computing $V^{\mathsf{SAFE}}_{m}\left(\mathcal{H}_{-}\right)$ directly, computing the upper bound of $mV^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$ also gives a tight bound.
D.1.1. Construction of $\mathcal{H}_{-}$.
@@
Table 3: Bid history achieving tight lower bound for $m=2$. Each round in the same partition has identical competing bid profile. Total number of units in each auction is $K=N^{3}+1$.
@@
D.1.2. Computing $\mathbf{b}_{m}^{\mathsf{OPT}}\left(\mathcal{H}_{-}\right)$.
@@
D.1.3. Computing $V^{\mathsf{SAFE}}_{1}\left(\mathcal{H}_{-}\right)$